OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 8x - y = 31 5x + 3y = 23

OpenStudy (anonymous):

use the first equation to get y in terms of x. in other words, get y by itself and then substitute it in the second equation.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

8x-y=31 ------> y = 8x-31 substitute that in the second equation and post what you get.

OpenStudy (anonymous):

A(4,1)

OpenStudy (anonymous):

8x-y=31 -- (*3)--> 24x-3y=93 (+) 5x +3y=23 (1) 29x=116 => x=4 (1)--x=4---> 5*4+3y=23 => 3y=23-20=> y=1 (4,1)